OpenStudy (anonymous):

Simplify the square root 98 times x to the fifth power.

OpenStudy (anonymous):

\[98=49\times 2\implies \sqrt{98}=\sqrt{49\times 2}=\sqrt{49}\sqrt{2}=7\sqrt{2}\] is a start